OpenStudy (anonymous):

Write an equation for a line parallel to x+15=-5y passing through (0,4)

OpenStudy (anonymous):

first put equation into y=mx+b then use point slope form y-y1=m(x-x1) m=slope

OpenStudy (anonymous):

0=x1 4=y1

OpenStudy (anonymous):

then the expression will look like y-4+m(x-0)

OpenStudy (anonymous):

@twinkieexpress ?

OpenStudy (anonymous):

slope will be -1/5 as it is parallel to the rrquired eqn.so slopes of oth equations wll be same.

OpenStudy (anonymous):

u r given a point (0,4) this will be x1 and y1 in the eqn, (y-y1)=slope(x-x1) put the values and try it

OpenStudy (anonymous):

How did you figured out the slope was -1/5?

OpenStudy (anonymous):

First substitute the value to make y positive: y=-0.2x-3 then use the gradient equation: (y-y1)=m(x-x1) where m is the slope and the point is (x1,y1) and you will get this: y-4=-0.2(x-0) and after simplifying you get: y=-0.2x+4

OpenStudy (anonymous):

because the slope on the x value is -1/5 when the coefficient of y is 1

OpenStudy (anonymous):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

what's the slope of x+15=-5y

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

you need to get the equation into y = mx+b form first

OpenStudy (anonymous):

is it y=1/5x+15

jimthompson5910 (jim_thompson5910):

close, but no

OpenStudy (anonymous):

umm can you help me with it, y intercept form was a sucky thing for me

jimthompson5910 (jim_thompson5910):

start with x+15=-5y and divide both sides by -5 to isolate y

jimthompson5910 (jim_thompson5910):

what you have is close, but it's a bit off

OpenStudy (anonymous):

umm 7=x-3

jimthompson5910 (jim_thompson5910):

hmm how are you getting that?

OpenStudy (anonymous):

umm im trying to divide andthen divide both sides by -5

jimthompson5910 (jim_thompson5910):

another way to think about it: divide every term by -5 to isolate y

OpenStudy (anonymous):

-1/5-3=y

jimthompson5910 (jim_thompson5910):

getting closer

OpenStudy (anonymous):

ugh

jimthompson5910 (jim_thompson5910):

you lost x

OpenStudy (anonymous):

y=1/5x-3

jimthompson5910 (jim_thompson5910):

it should be -1/5x - 3 = y

jimthompson5910 (jim_thompson5910):

much better

jimthompson5910 (jim_thompson5910):

oh sry

jimthompson5910 (jim_thompson5910):

it should be y = -1/5x - 3 but you pretty much have it (ignoring the typo)

OpenStudy (anonymous):

and whats my next step?

jimthompson5910 (jim_thompson5910):

so what is the slope of y = -1/5x - 3

OpenStudy (anonymous):

-1/5

jimthompson5910 (jim_thompson5910):

so the slope of any parallel line to this one is also -1/5

jimthompson5910 (jim_thompson5910):

since parallel lines have equal slopes

jimthompson5910 (jim_thompson5910):

the given point is (0,4) so x = 0 and y = 4 -------------------- we know that m = -1/5, x = 0, y = 4 plug these 3 pieces of info into y = mx + b and solve for b

OpenStudy (anonymous):

y=1/5x+4

jimthompson5910 (jim_thompson5910):

the slope is negative, so it should be y = -1/5x + 4